# Can Graph Learning Improve Task Planning?

### v1

@@ 4.2 A Training-free GNN-based Approach
We introduce GNNs into the task retrieval stage by formulating it as a node classification problem. The node embedding is the embedding of this node’s description. For example, the node description in HuggingFace can be “Translation. Translation is the task of converting text from one language to another”. Given that the node description is usually simple, node embedding is obtained by e5-355M [44] for efficiency. Then we use the parameter-free GNN (i.e., SGC [47]) to refine the node embeddings. The update is given by $\bm{H}^{(k)}=\alpha\bm{H}^{(k-1)}+(1-\alpha)\hat{\bm{A}}\bm{H}^{(k-1)}$, where $\bm{H}^{(k)}$ is the embedding after $k$-layer SGC, $\bm{H}^{(0)}$ is the initial node feature, $\hat{\bm{A}}$ is the normalized adjacency matrix of the task graph, and $\alpha$ is a hyperparameter.
When selecting the task node for a step $s_{i}$, e5-355M is used to embed the LLMs’ generated step as $\bm{x}_{i}^{\text{step}}$. Given a sequence of previously selected task nodes $\{v_{1},\cdots,v_{i-1}\}$, the next node $v_{i}$ is chosen according to $v_{i}=\text{argmax}_{v\in\mathcal{N}(v_{i-1})}\;\langle\bm{h}_{v},\bm{x}_{i}^{\text{step}}\rangle$, where $\bm{h}_{v}$ is the node embedding. Particularly, $v_{1}$ can be selected from the whole graph. The one-by-one node selection approach is inspired by the traditional research of GNNs for decision-making problems like combinatorial optimization [16].
@@ 4.3 A Training-required GNN-based Approach
Data Preparation: We assume that each entry in the task planning dataset comprises a user request, a sequence of decomposed steps, and the corresponding ground-truth tasks, denoted as $(\text{request},\{s_{1},\ldots,s_{n}\},\{v_{1},\ldots,v_{n}\})$. If the dataset does not adhere to this format, we reformat it accordingly using GPT-4, with details provided in Appendix E.2. It is important to note that there is a one-to-one correspondence between the steps and tasks in the dataset. The node classification dataset is represented as $\{s_{i},v_{i}\}_{i=1}^{n}$, where $s_{i}$ is a step described in natural language, and $v_{i}$ is its corresponding task.
Training Loss: Inspired by GNN for recommendation [12], we adopt the Bayesian Personalized Ranking (BPR) loss [27] designed for recommendation with binary rankings. The loss function is given by
$\ell=\sum_{(\bm{x}^{\text{step}},v,v^{\prime})}-\log\;\sigma(\langle\bm{h}_{v},\bm{x}^{\text{step}}\rangle-\langle\bm{h}_{v^{\prime}},\bm{x}^{\text{step}}\rangle),$
where $\bm{x}^{\text{step}}$ represents the embedding of the step’s textual description generated by e5-355M, $v$ is ground-truth task, and $v^{\prime}$ is a negative task. We select negative tasks that are textually similar to the positive task, and for computational efficiency, we limit our selection to $2$ negative tasks per positive task. The trainable parameters may merely include GNNs or both GNNs and e5-355M.
@@ 5.1 Experimental Setup
Evaluation: For the datasets from TaskBench, we split $3000$ samples for training and $500$ samples for testing, each containing an invocation path with at least two tasks. For the TMDB dataset, we first filter to include the samples with two or more invoked tasks, and then randomly select a sample served as the in-context learning example. The remaining $94$ samples are designated for testing. We adopt the evaluation metric in TaskBench [30] and HuggingGPT [31], i.e., Node F1-Score (n-F1) and Link F1-Score (l-F1), which measure the accuracy of invoked tasks and invoked dependencies, respectively. We also measure the token consumption (# tok) as the efficiency metric.
@@
Choices of GNNs: To comprehensively investigate the effectiveness of different graph learning methods in enhancing task planning, we consider a wide range of graph neural networks, including SGC [47], GCN [18], GAT [39], GraphSAGE [11], GIN [49], and TransformerConv [32].
@@ 5.2 Performance of the Training-free Approach
We compare the performance across three training-free methods: (1) LLM’s Direct Inference is introduced in Section 4.1. (2) GraphSearch [23, 36, 22] leverages the classic graph search method to generate the candidate nodes and use LLMs to give a score to the selection. Given a step, GreedySearch consistently selects the node with the highest score and adjacent to the previous task node; AdaptiveSearch selects the nodes with scores above a fixed threshold, adjusting the breadth of the search space in an adaptive mode; BeamSearch retains the $k$ nodes with highest scores. (3) SGC [47]: Employ a training-free SGC for task retrieval based on decomposed task steps. Illustrations of these methods are shown in Figures 5 and 6 in Appendix. Table 1 shows both the overall performance and token consumption costs.
@@
Compared with direct inference, integrating SGC consistently improves performance, underscoring the effectiveness of the proposed method. GraphSearch-type methods rely on beam search to identify paths and employ LLMs for evaluation, where longer processing times generally lead to better outcomes. Notably, our proposed method achieves comparable or superior performance to BeamSearch while requiring $5$-$10$ times fewer tokens (Table 1) and inference time (Table 6). The case studies are provided in Appendix F.4. However, we observed only marginal improvements with GPT-4-turbo. A unique feature of GPT-4-turbo is its ability to manage ChatGPT-plugins, and it may have been specially trained on task planning datasets. In addition, the language model used for feature extraction in SGC is e5-355M, which may not be sufficiently powerful to effectively analyze GPT-4’s output. A detailed diagnostic analysis of cases involving GPT-4 is provided in Figure 11.
@@ 5.3 Performance of the Training-based Approaches
Settings: We further explore the efficacy of training-based GNNs in three TaskBench datasets. The TMDB dataset is excluded due to its limited sample size. Throughout our experiments, we trained a spectrum of GNN variants, both with and without co-training the small LM, whose role is to generate node embeddings derived from task names and descriptions. Owing to space constraints, we only show the performance of GraphSAGE in the main text, relegating a detailed comparison of all the situations to Table 8 and Table 9.
Observations: From Table 2, we observe a significant improvement in performance when employing a training-based GraphSAGE approach over the training-free method. However, the co-training of GNNs with e5-355M does not yield a marked improvement, suggesting that message passing is the crucial element for enhancing performance. Further analysis across a broad spectrum of GNNs (as shown in Table 8 and Table 9) reveals that powerful GNNs, such as GINs, perform similarly to networks perceived as less complex, like GCNs, and even underperform compared to GraphSAGE. This pattern indicates that the task’s challenge may not lie in the expressiveness of the models but rather in their ability to generalize.
Efficiency: The training cost is remarkably low because we use e5-355M [44] as the text embedding model for GNNs. If the trainable parameters are limited to the GNNs alone, training typically concludes within just 3 minutes. Furthermore, the training duration extends to only 15 minutes when GNNs are jointly trained with e5-355M model. This efficiency stands in stark contrast to the 10-20 hours required for tuning open-sourced LLMs. The details of the training time are given in Table 7.
Table 2: Performance of training-based GNN. The performance of all other GNNs are given in Table 8 and Table 9 in the Appendix.
@@
Baichuan2-13B Direct 45.85 19.00 47.57 4.08 33.45 9.52
GraphSAGE 59.32 34.36 56.15 31.60 65.18 40.49
GraphSAGE (co-train) 59.76 35.59 57.97 33.29 63.21 38.10
Vicuna-13B Direct 50.46 21.27 53.57 23.19 73.70 45.80
GraphSAGE 61.86 35.68 63.71 39.88 86.07 67.63
GraphSAGE (co-train) 62.82 37.04 65.89 42.18 84.23 65.44
CodeLlama-7B Direct 58.06 29.39 59.44 30.83 84.12 62.89
GraphSAGE 66.67 43.03 67.97 46.31 88.53 72.02
GraphSAGE (co-train) 67.19 42.94 70.00 48.28 87.81 70.20
Mistral-7B Direct 60.60 30.23 69.83 39.85 84.26 53.63
GraphSAGE 68.12 43.09 75.51 52.94 87.51 66.57
GraphSAGE (co-train) 67.61 43.14 76.96 55.46 87.61 66.75
CodeLlama-13B Direct 57.55 28.88 68.57 41.79 91.20 76.07
GraphSAGE 67.30 42.41 74.93 54.52 93.84 80.38
GraphSAGE (co-train) 68.92 44.85 76.28 55.41 93.30 79.51
GPT-3.5-turbo Direct 73.85 45.73 82.85 62.07 96.09 83.65
GraphSAGE 77.90 52.68 85.29 65.80 96.43 86.26
GraphSAGE (co-train) 77.87 53.04 85.51 66.56 96.34 86.09
GPT-4-turbo Direct 77.60 52.18 88.29 69.38 97.36 84.58
GraphSAGE 78.76 52.53 88.63 69.65 97.34 85.67
GraphSAGE (co-train) 78.49 52.62 88.86 70.25 97.42 85.80
@@
Figure 3: Orthogonal Effectiveness of GNN on either different prompt templates (on HuggingFace dataset) or fine-tuned LLMs.
@@ 5.4 Improved Prompts and Fine-tuned LLMs
In this subsection, we show that the proposed method is orthogonal to two dominant methods, i.e., the prompt engineering and finetuning.
Orthogonal to Improved Prompts: We investigate GNN’s effectiveness when applied to improved prompt templates, i.e., strategically designed prompts that enhance the task planning abilities of LLMs. Specifically, we consider two types of prompts: (1) In-context Learning with Increased Examples [30] During main experiments, we maintain the consistent 1-shot in-context learning example for LLM’s direct inference. To realize further improvements, we increase the number of examples to $2$, and results under this setting are denoted as “2-shot Prompt”; (2) Plan like a Graph (PlaG) [20] We adopt the prompt in [20] to encourage LLM to think and plan in a graph-like manner, we convert the entire task graph into plain text and then integrate PlaG instructions to enhance LLM’s task planning. Results under this prompt template are denoted as “PlaG Prompt”.
From the results shown in Figure 3(a), where we apply three different prompts to CodeLlama-13B and Mistral-7B on HuggingFace, it is clear that applying a GNN to improved prompts, where task steps are more concisely decomposed and predictions are more accurate, can also boost performance.
Orthogonal to LLMs’ Fine-tuning: To explore whether our framework maintains effectiveness on fine-tuned LLMs, which have acquired dataset-specific task planning capabilities, we conduct further experiments. For each dataset, we use LoRA [13] to fine-tune two LLMs of different parameter scales, including CodeLlama-7B and Vicuna-13B, based on the same training data as GNNs. Details of fine-tuning process are provided in Appendix D.4. The finetuned model is named as “FT-CodeLlama” and “FT-Vicuna” in Figure 3(b).
The results depicted in Figure 3(b) demonstrate that fine-tuning markedly enhances the task-planning capabilities of LLMs. Furthermore, applying GraphSAGE to the decomposed tasks of LLMs further improves the accuracy of the task planning.
6 Related Works
Due to space limitations, this section focuses on related works of task planning, while discussions of other relevant areas are deferred to Appendix A. The existing studies of task planning approaches can be categorized into several directions, including task decomposition, multi-plan selection, the use of external planners, reflection, and memory-aided planning [14]. Task decomposition methods, such as the chain-of-thought approach [45], employ the divide-and-conquer strategy, utilizing LLMs for both task decomposition and sub-task planning. The application of this method to task planning is detailed in Section 4.1 and is referred to as “Direct” in the baseline comparison. Multi-plan selection strategies, exemplified by the tree-of-thought [53] and graph-of-thought [2], leverage search-based methods to generate plans. Subsequently, LLMs evaluate these plans to select the most effective one. The GraphSearch methods used in our baselines fall into this category. External planner approaches [21] formalize tasks using the Planning Domain Definition Language (PDDL) and employ classic solvers to address the planning problem, which is closely related to the proposed approaches in this paper. The task planning considered here is more flexible and cannot be directly translated into formal languages. However, we demonstrate that GNNs can serve as an effective external planner. Reflection-based methods [33] focus on reflecting upon experiences to refine the plan, while memory-aided planning approaches [55] utilize external experiences, such as those from search engines. These approaches are deployed in interactive environments and orthogonal to this paper.

### v2

@@ 4.2 A Training-free GNN-based Approach
As we discussed in Section 2.2, the current solution to task planning involves two stages. The first stage requires the ability to understand users’ requests in natural language and break them down into concrete instructions, which is the unique ability of LLMs. The second stage is to select a path on the task graph, where each node corresponds to a decomposed step. Thus, we can integrate GNNs in this stage. The illustration of our framework is shown in Figure 7 in Appendix.
For each decomposed step outputted by the first stage, we use a GNN to select a corresponding node within the task graph. Suppose we are selecting the node for the $i$-th decomposed step. First, we utilize a small pre-trained language model, e5-355M [61], to embed the $i$-th decomposed step. The resulting embedding is denoted as $\bm{x}_{i}^{\text{step}}$. Second, for the task graph, we first use the same pre-trained language model e5-355M to convert each node’s description into embeddings, denoted as the node feature $\bm{h}^{0}_{v}$, where the superscript indicates the layer and the subscript represents the node. Then we adopt a $K$-layer SGC [64] to compute the final node embeddings, resulting in $\bm{h}_{v}=\bm{h}_{v}^{(K)}$. Given a sequence of previously selected task nodes $\{v_{1},\cdots,v_{i-1}\}$, the next node $v_{i}$ is chosen according to $v_{i}=\text{argmax}_{v\in\mathcal{N}(v_{i-1})}\;\langle\bm{h}_{v},\bm{x}_{i}^{\text{step}}\rangle$, where $\bm{h}_{v}$ is the final node embedding, and $\mathcal{N}(v_{i-1})$ denotes the neighbors of node $v_{i-1}$ in the task graph. Particularly, $v_{1}$ can be selected from the whole graph.
As e5-355M is pre-trained and SGC is parameter-free, the proposed method requires no additional training and can be effectively applied in a zero-shot manner.
@@ 4.3 A Training-required GNN-based Approach
Data Preparation: We assume that each entry in the task planning dataset comprises a user request, a sequence of decomposed steps, and the corresponding ground-truth tasks, denoted as $(\text{request},\{s_{1},\ldots,s_{n}\},\{v_{1},\ldots,v_{n}\})$. If the dataset does not adhere to this format, we reformat it accordingly using GPT-4, with details provided in Appendix C.2. It is important to note that there is a one-to-one correspondence between the steps and tasks in the dataset. Therefore, the training dataset can be represented as $\{(s_{i},v_{i})\}_{i=1}^{n}$, where $s_{i}$ is a step described in natural language, and $v_{i}$ is its corresponding invoked task.
Training Loss: The problem in the dataset can be viewed as a binary ranking problem, where the labeled node is $1$ and the other nodes are 00. Therefore, we adopt the Bayesian Personalized Ranking (BPR) loss [40] designed for recommendation with binary rankings. The loss function is given by $\ell=\sum_{(\bm{x}^{\text{step}},v,v^{\prime})}-\log\;\sigma(\langle\bm{h}_{v},\bm{x}^{\text{step}}\rangle-\langle\bm{h}_{v^{\prime}},\bm{x}^{\text{step}}\rangle)$, where $\bm{x}^{\text{step}}$ represents the embedding of the step’s textual description generated by e5-355M, $v$ is the ground-truth task, and $v^{\prime}$ is a negative task. We select negative tasks that are textually similar to the positive task, and for computational efficiency, we limit our selection to $2$ negative tasks per positive task. The trainable parameters may merely include GNNs or both GNNs and e5-355M with illustrative configurations shown in Figure 8 in Appendix.
@@ 5.1 Experimental Setup
Evaluation: For the datasets from TaskBench, we split $3000$ samples for training and $500$ samples for testing, each containing an invocation path with at least two tasks. For the TMDB dataset, we first filter to include the samples with two or more invoked tasks, and then randomly select a sample served as the in-context learning example. The remaining $94$ samples are designated for testing. We adopt the evaluation metric in TaskBench [44] and HuggingGPT [45], i.e., Node F1-Score (n-F1) and Link F1-Score (l-F1), which measure the accuracy of invoked tasks and invoked dependencies, respectively. Besides, the Accuracy (Acc) can measure the success rate from task level. We also measure the token consumption (# tok) as the efficiency metric.
@@
Choices of GNNs: To comprehensively investigate the effectiveness of different graph learning methods for task planning, we consider a wide range of graph neural networks, including SGC [64], GCN [25], GAT [56], GraphSAGE [15], GIN [66], and Graph Transformers [46].
@@ 5.2 Performance of the Training-free Approach
We compare the performance across three training-free methods: (1) LLM’s Direct Inference is introduced in Section 2.2. (2) GraphSearch [32, 51, 31] leverages the classic graph search method to generate the candidate nodes and use LLMs to give a score for node selection. Given a step, GreedySearch consistently selects the node with the highest score and adjacent to the previous task node; AdaptiveSearch selects the nodes with scores above a fixed threshold, adjusting the breadth of the search space in an adaptive mode; BeamSearch retains the $k$ nodes with highest scores. (3) SGC [64] employs a training-free SGC for task retrieval based on decomposed task steps. The details of baselines are given in Section E.1 and illustrated in Figure 6. Table 1 shows both the overall performance and token consumption costs, with results of Accuracy (Acc) moved to Table 9 in Appendix.
@@
Compared with direct inference, integrating an SGC consistently improves performance, underscoring the effectiveness of the proposed method. GraphSearch-type methods rely on beam search to identify paths and employ LLMs for evaluation, where longer processing times generally lead to better outcomes. Notably, our proposed method achieves comparable or superior performance (Table 1 and Table 9) to BeamSearch while requiring $5$-$10$ times fewer tokens (Table 1) and inference time (Table 10). The case studies are provided in Appendix H. However, we observed only marginal improvements with GPT-4-turbo. A unique feature of GPT-4-turbo is its ability to manage ChatGPT-plugins, and it may have been specially trained on task planning datasets. In addition, the pre-trained language model used for feature extraction in SGC is e5-355M, which may not be sufficiently powerful to effectively analyze GPT-4’s output. A detailed diagnostic analysis of cases involving GPT-4 is provided in Figure 11 in Appendix.
@@ 5.3 Performance of the Training-based Approaches
Settings: We further explore the efficacy of training-based GNNs in three TaskBench datasets. The TMDB dataset is excluded due to its limited sample size. Throughout our experiments, we trained a spectrum of GNN variants, both with and without co-training the small LM (i.e., e5-355M), whose role is to generate node embeddings derived from task names and descriptions. Owing to space constraints, we only show the performance of GraphSAGE in the main text, relegating a detailed comparison of all the situations to Table 11 and Table 12 in Appendix.
[…]
Observations: From Table 2, we observe a significant improvement in performance when employing a training-based GraphSAGE approach over the training-free method. However, the co-training of GNNs with e5-355M does not yield a marked improvement, suggesting that message passing is the crucial element for enhancing performance. Further analysis across a broad spectrum of GNNs (as shown in Table 11 and Table 12) reveals that powerful GNNs, such as GINs, perform similarly to networks perceived as less complex, like GCNs, and even underperform compared to GraphSAGE. This pattern indicates that the task’s challenge may not lie in the expressiveness of the models but rather in their ability to generalize. Regarding baselines, TAPE is unsuitable for task planning as its classification approach simplifies task planning, overlooking task dependencies. While GraphToken demonstrates superior performance over LLMs’ direct inference, we have noted instances of minor hallucination. This observation suggests that GraphToken’s understanding of the task graph is not yet perfect. In addition, GraphToken is limited to open-sourced LLMs. The proposed approaches also greatly improve the parameter prediction performance, as given in Appendix G (e.g. $9\%$ improvement to to GPT-3.5 and $3\%$ improvement to GPT-4).
Efficiency: The details of the training time are given in Table 13. The training cost is remarkably low because we use e5-355M [61] as the text embedding model for GNNs. If the trainable parameters are limited to the GNNs alone, training typically concludes within just 3 minutes. Furthermore, the training duration extends to only 15 minutes when GNNs are jointly trained with e5-355M model. This efficiency stands in stark contrast to the 10-20 hours required for tuning open-sourced LLMs.
[…]
@@
LLM Method n-F1 $\uparrow$ l-F1 $\uparrow$ Acc $\uparrow$ n-F1 $\uparrow$ l-F1 $\uparrow$ Acc $\uparrow$ n-F1 $\uparrow$ l-F1 $\uparrow$ Acc $\uparrow$
Vicuna-13B Direct 50.46 21.27 8.72 53.57 23.19 11.20 73.70 45.80 24.43
TAPE 59.47 NA 5.07 54.97 NA 2.07 73.26 NA 12.50
GraphToken 63.37 31.54 15.61 65.40 36.38 19.87 81.65 48.29 43.37
GraphSAGE 61.86 35.68 20.08 63.71 39.88 21.37 86.07 67.63 48.64
[…]
Mistral-7B Direct 60.60 30.23 16.36 69.83 39.85 25.05 84.26 53.63 44.52
TAPE 61.82 NA 6.13 58.92 NA 3.29 76.40 NA 16.44
GraphToken 64.42 32.04 18.60 72.31 42.60 30.31 86.82 57.06 53.99
GraphSAGE 68.12 43.09 25.77 75.51 52.94 34.29 87.51 66.57 52.74
[…]
CodeLlama-13B Direct 57.55 28.88 14.29 68.57 41.79 24.10 91.20 76.07 66.40
TAPE 64.03 NA 8.05 58.27 NA 2.01 77.74 NA 17.37
GraphToken 62.15 32.55 20.08 74.57 47.60 35.06 92.50 73.57 69.42
GraphSAGE 67.30 42.41 26.56 74.93 54.52 38.55 93.84 80.38 73.60
[…]
Direct 73.85 45.73 28.95 82.85 62.07 47.96 96.09 83.65 81.30
TAPE 68.00 NA 8.83 62.43 NA 3.87 70.67 NA 8.92
GPT-3.5-turbo GraphSAGE 77.90 52.68 35.11 85.29 65.80 53.55 96.43 86.26 83.13
[…]
Direct 77.60 52.18 33.68 88.29 69.38 60.56 97.36 84.58 86.77
TAPE 68.82 NA 9.50 63.94 NA 4.02 71.51 NA 9.40
GPT-4-turbo GraphSAGE 78.76 52.53 34.09 88.63 69.65 60.36 97.34 85.67 86.97
[…]
Table 3: Performance Comparison of SGC and GraphSAGE on the UltraTool Benchmark [19]. Integrating GNNs can lead to more significant improvements in LLMs on larger task graphs.
0-shot 1-shot
LLM Method n-F1 $\uparrow$ l-F1 $\uparrow$ Acc $\uparrow$ # Tok $\downarrow$ n-F1 $\uparrow$ l-F1 $\uparrow$ Acc $\uparrow$ # Tok $\downarrow$
Direct 38.88 16.42 13.58 10,535 57.64 30.44 26.25 10,737
BeamSearch 49.71 22.51 17.08 26,008 64.93 36.23 33.47 23,023
SGC 61.07 37.61 25.31 10,456 71.64 44.00 39.68 10,658
CodeLlama-13B GraphSAGE 63.78 39.91 27.98 10,456 72.81 45.26 43.49 10,658
Direct 54.35 21.35 18.33 8,462 63.58 30.85 25.00 8,614
BeamSearch 55.40 28.02 19.76 21,979 63.41 34.05 26.28 20,813
SGC 59.80 37.82 25.87 8,352 64.96 37.96 29.70 8,504
GPT-3.5-turbo GraphSAGE 63.97 42.26 30.35 8,352 70.49 47.79 39.74 8,504
Direct 68.63 40.01 27.20 8,513 69.54 41.79 28.17 8,693
BeamSearch 71.29 43.99 30.40 18,793 71.99 44.54 31.62 20,515
SGC 70.87 44.01 31.60 8,346 70.46 44.82 33.00 8,504
GPT-4-turbo GraphSAGE 70.67 43.83 34.40 8,346 70.75 47.68 37.22 8,504
5.4 Scaling to Large Task Graphs
To demonstrate the scalability of our method to larger task graphs, we conducted a supplementary experiment on the newly released planning benchmark UltraTool [19], which features a relatively large task graph with $260$ nodes. Details on processing this dataset are provided in Appendix C.3. We present a performance comparison of GNN models (training-free SGC and training-required GraphSAGE) against strong baselines like BeamSearch in Table 3. Among the metrics, accuracy (Acc) is calculated based on whether the predicted tasks match the ground-truth tasks, measuring the success rate at each case level. In such conditions, integrating a GNN significantly enhances performance and mitigates planning failures, e.g., GPT-4-turbo undergoes a 9.05% accuracy improvement with the introduction of GraphSAGE.
The results indicate that (1) LLMs’ performance is vulnerable to the scale of task graphs; (2) The performance gain of the proposed method increases with a larger task graph.
@@ 5.4 Scaling to Large Task Graphs
Figure 3: Orthogonal Effectiveness to both Improved Prompts and Fine-tuned LLMs
@@ 5.5 Improved Prompts and Fine-tuned LLMs
In this subsection, we show that the proposed method is orthogonal to two dominant methods, i.e., prompt engineering and fine-tuning.
Orthogonal to Improved Prompts: We investigate GNN’s effectiveness when applied to improved prompt templates, i.e., strategically designed prompts that enhance the task planning abilities of LLMs. Specifically, we consider two types of prompts: (1) In-context Learning with Increased Examples [44] During main experiments, we maintain the consistent 1-shot in-context learning example for LLM’s direct inference. To realize further improvements, we increase the number of examples to $2$, and results under this setting are denoted as “2-shot Prompt”; (2) Plan like a Graph (PlaG) [29] We adopt the prompt in [29] to encourage LLM to think and plan in a graph-like manner. Specifically, we convert the entire task graph into plain text and then integrate PlaG instructions to enhance LLM’s planning. Results under this prompt template are denoted as “PlaG Prompt”.
From the results shown in Figure 3(a), where we apply three different prompts to CodeLlama-13B and Mistral-7B on HuggingFace, it is clear that applying GraphSAGE to improved prompts, where task steps are more concisely decomposed and predictions are more accurate, can also boost performance.
Orthogonal to LLMs’ Fine-tuning: To explore whether our framework maintains effectiveness on fine-tuned LLMs, which have acquired dataset-specific task planning capabilities, we conduct further experiments. For each dataset, we use LoRA [18] to fine-tune two LLMs of different parameter scales, including CodeLlama-7B and Vicuna-13B, based on the same training data as GNNs. Details of fine-tuning process are provided in Appendix F.3. The finetuned model is named as “FT-CodeLlama” and “FT-Vicuna” in Figure 3(b).
The results depicted in Figure 3(b) demonstrate that fine-tuning markedly enhances the task-planning capabilities of LLMs. Furthermore, applying GraphSAGE to the decomposed tasks of LLMs further improves the accuracy of task planning.